Can someone please help me with this equation
19h - 7 = 23 + 14h

Answer by stanbon(75887)

Can someone please help me with this equation
19h - 7 = 23 + 14h
19h-14h = 23+7
5h = 30
h = 6
-------------
Cheers,
Stan H.